If utility function is given by U(C,L) = (C-350).(L-10) calculate reservation
wage if total leisure hours is 110 and income without work is $500. (Where C
= consumption, and L = leisure)
1
Expert's answer
2020-03-02T08:12:30-0500

The reservation wage is defined as MRS when working no hours.

"wRES = \frac{MUL}{MUC} = \frac{C - 350}{L - 10} = \frac{500 - 350}{110 - 10} = 150/100 = \$1.5."